Parents set up a trust fund for their children last year. Part of it is 8% simple interest per year while the rest of it is earning 6% simple interest per year. They placed $4000 in the trust fund. In one year the trust fund has earned $280. How much did they invest at each interest rate?

Answer by richwmiller(17219)

x+y=4000,
0.08*x+0.06*y=280
x=4000-y
0.08*(4000-y)+0.06*y=280
320-0.08y+0.06*y=280
-0.02*y=-40
y=2000 at 6%
x=4000-y
x=2000 at 8%
check
0.08*2000+0.06*2000=280
160+120=280
280=280
ok